OpenStudy (anonymous):

Solve for n -10+log3(n+3)=-10

OpenStudy (anonymous):

n=-8/3 -10+log3(n+3)=-10 log3(n+3)=0 log3(n+3)=log1 .....(log1=0) By comparing, 3(n+1)=1 n=(1/3)-3= -8/3

OpenStudy (anonymous):

I disagree with the above comment, though I could be wrong. I got n=-2 10+log3(n+3)=-10 log3(n+3)=0 log((n+3)^3)=0 <-- this step is because of the exponential log property (n+3)^3=1 n=-2 However, I could be wrong.

OpenStudy (anonymous):

The property is actually like this a(logn)= log(n^a) also log(ab)=loga+logb